OpenStudy (anonymous):

*Advance Algebra With Financial Applications* APR=(1+0.0155833333)^(12)(30)-1 ? APR 18.7% compounded monthly Grace Period 30 days Minimum Payment 3.5% of the remaining balance or $30, whichever is higher Pamela charged $925 to her credit card last month. What is the first minimum payment she must make? $92.50 $172.98 $32.38 $30.00

OpenStudy (anonymous):

Is this the correct formula ? APR=(1+0.0155833333)^(12)(30)-1

OpenStudy (anonymous):

APY=(1+0.0155833333)^(12)-1 APY=(1.0155833333)^(12)-1 APY=1.2039 - 1 APY=0.2039 =20.39 % ?

OpenStudy (amistre64):

\[(1+\frac{apr}{12})^{12}=1+apy\] \[(1+\frac{apr}{12})^{12}-1=apy\]

OpenStudy (amistre64):

but no, thats not what the question is asking for

OpenStudy (amistre64):

B(1+apr/12) = ______

OpenStudy (anonymous):

(1+0.187/12)=1 + 0.2039 ?

OpenStudy (amistre64):

i forgot to factor in the grace period, but even then the amount differs by ony cents

OpenStudy (amistre64):

3.5% of 925, or $30 which is greater?

OpenStudy (anonymous):

$30

OpenStudy (amistre64):

925(.052) = 32.38 30.00 ..... 32.00 ^^^^^

OpenStudy (anonymous):

where .052 from ?

OpenStudy (amistre64):

i tried to fix my .032 into .035 but editied the wrong number :/ even if she went over the grace period: 925(1+.187/12)(.035) = 32.88

OpenStudy (anonymous):

lol i feel like an idiot that all I had to do ?

OpenStudy (amistre64):

thats all i read it as ...

OpenStudy (anonymous):

Thank you soo much ! :D